2 kg particle starts at the origin and moves along x axis .the net force measured at intervals of 1m are 28,30,40,48,50,64.5,13.5,12.2,40.8,27,23 Newtown .find total work done
1
Expert's answer
2020-10-05T10:56:20-0400
"W=\sum (Fx)=\sum F \cdot Nd\\\sum F=28+0+40+48+50+64.5\\+13.5+12.2+40.8+27+23=347\ N"

"W=347(11)(1)=3817\ J"